HUBBLE SPACE TELESCOPE DAILY REPORT #956 PERIOD COVERED: 0800 EST Daily Status Report as of 235/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of four observations were received. The observations were reported to be normal. 1.2 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make two internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. Both observations were executed as scheduled, and no problems were reported. 1.3 Completed FGS 4935 (Parallaxes of Hyades Cluster Members Cycle 3 High) The Fine Guidance Sensors were used to make astrometric observations of star VA-310 which is located in the Hyades cluster to obtain a definitive calibration of its population I zero-age main sequence members. There was real-time telemetry for the entire set of observations. There were two recentering events which occurred during the observations, otherwise no problems occurred. 1.4 Completed Two Sets of PC 4940 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) The Planetary Camera was used to make three images of the planetary nebula SMC-SMP-22 which is located in the Small Magellanic Cloud. These observations will provide information on the morphology, size, and dynamical age of the object. All three observations were displayed, and they were reported to look okay. 1.5 Completed PC 4904 (The Nuclei of Nearby S0 and Spiral Galaxies -III A PC Snapshot Image) The Planetary Camera was used to make one observations of the low-red-shift galaxy NGC-2640 to study its surface brightness profile, identify sub-arc-second scale condensations, find point sources, and look for gasious disk and dust lanes. The observation was executed without problems. The data was received and displayed. No problems were reported. 1.6 Completed PC 4785 (WF/PC Cycle 2 Calibration: Interchip Photometric Calibration) The Wide Field Camera and Planetary Cameras were used to make 25 calibration images of standard star HD+75D325 to help calibrate the detector quantum efficiency of each of the Wide Field and Planetary Camera chips. All 25 observations were received and displayed. The observations were reported to look normal. 1.7 Completed GHRS/2 4100 (Highly Evolved Stars of Low Mass) The High Resolution Spectrograph was used to make observations of standard star BD+28D4211. This is a hot sub-dwarf standard star used by both IUE and HST. A total of nine observations were taken. recentering evens were reported to have occurred during some of the observations. Despite the recentering events, the spectral observations were reported to look good. 1.8 Completed PC 3282 (Binaries in Globular Clusters (FOS 37): Cycle 2 Observations) The Planetary Camera was used to search for optical counterparts of X-Ray sources which are located in globular cluster NGC-1851. The candidate objects are expected to be very faint. A total of six observations were taken. There was some sort of data transmutation problems reported on both the DOC and DCF during a data dump of the science tape recorder during one of the observations. Other than that, the exposures were reported to be good. 1.9 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.10 Completed Three Sets of PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) The Wide Field Camera was used to make one bias and one dark calibration images. These data are required for routine calibration of the instrument and for data reduction purposes. Both of the observations were reported to be normal calibration images. 1.11 Completed FOS/BL/RD 4976 (UV Rotational Light Curs for Pluto, and Charon's UV Spectrum) The Faint object Spectrograph was used to make a TALED exposure subsequent to other spectral observations of the planet Pluto and its moon Charon. The TALED observation was received in real time. It was reported to look okay. After a corrective SIP maneuver was uplinked to the spacecraft, spectral observations were made of the planet Pluto. The observations were executed in fine lock, and received in real time. The spectral observations were normal. 1.12 Completed FGS 5064 (Diagnostic of Early Stellar Evolution: Search for Pre-main Sequence Stars) Fine Guidance Sensor #3 was used to make observations of pre-main sequence star 20160 (aka, BP Tau) to determine whether it has a companion. The observations were executed as scheduled, and no problems were reported. 1.13 Completed PC 4308 (Snapshot Survey for Companions to Planetary Nebulae Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula ABELL-48. It is possible that the central star in a double star. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.14 Completed FOC 5028 (FOC Objective Prisms - Wave Length Calibrations Verification) The Faint Object Camera (f/96) was used to make some direct images and prism observations of symbiotic star HM-SGE. One set of observations were wavelenght calibrations, and an other set were be a comparison images. The observations were executed as scheduled, and no problems were reported. 1.15 Completed Two Sets of PC 4738 (WF/PC Cycle 3 Calibration: UV Flats 1) The WFPC was used to make one image of the Earth to provide flat field images for calibration of the ultraviolet observations. The image was received and displayed. The observation was received and displayed. No anomalies were reported. 1.16 Completed Two Sets of FGS 4892 (Duplicity Among Hyades Stars Cycle 3 High) The FGSs were used to make astrometric observations of the possible binary star VA-490 which is located in the Hyades open cluster. The OSS monitored the observation. One recentering event was reported to have occurred. Other than that no other anomalies were reported. 1.17 Completed FOS/BL 4396 (QSO Absorption systems Snapshot Survey - Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make two spectrographic observations of quasar 0355-4820 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The data was received and displayed. and no anomalies were reported. 1.18 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD182369 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 Scheduled Reacquisitions : 25 Successful : 25 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 37 Successful : 37 2.3 Significant HSTARs: None 2.4 Operations Note: The ground systems will start prepertations tonight for the Joint Integration Simulation #1 (JIS #1, Rendezvous to Solar Array Stow and EVA #1) which will be held from Tuesday Aug. 24 through Wednesday Aug. 25. The test should last approximately 37 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.2 complete FOC/96 5030 (F/96 Filter Photometric Calibrations) 3.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.5 Complete PC 5014 (WF/PC 3 Calibration: PC Flats Narrow. 2) 3.6 Complete HSP/UV1/UV2/POL 5082 (X-Ray binaries Part 2 - Cycle 3 High) 3.7 Complete HSP/UV2 4209 (Determination of the Masses and Densities of Pluto and Charon - cycle 3) 3.8 Complete PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) 3.9 Complete HSP/POL 4887 (X-Ray Binaries - AUG - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-